OpenStudy (yesgeorgie8):

For questions 1-8 verify that the given value is a solution for the trigonometric equation by showing that it makes the eqution true. PLEASE HELP ME OUT I NEED ANSWERS !!

OpenStudy (yesgeorgie8):

surjithayer @surjithayer

OpenStudy (anonymous):

i do not see any question here.

OpenStudy (yesgeorgie8):

For questions 1-8 verify that the given value is a solution for the trigonometric equation by showing that it makes the eqution true. PLEASE HELP ME OUT I NEED ANSWERS !!

OpenStudy (yesgeorgie8):

@surjithayer

OpenStudy (anonymous):

first two are correct.

OpenStudy (yesgeorgie8):

yes :)

OpenStudy (anonymous):

3.\[\csc x=2,\sin x=\frac{ 1 }{ 2 }=\sin \left( 2n \pi+\frac{ \pi }{ 6 } \right),\] where n is an integer. for n=1,what you get?

OpenStudy (yesgeorgie8):

\[\sin(2\pi+\pi/6)\]

OpenStudy (anonymous):

4. also \[\sin x=\sin \left( 2n \pi+\frac{ 5 \pi }{ 6 } \right),for ~n=1\] what you get?

OpenStudy (yesgeorgie8):

i don't really know how to solve or do these :/

OpenStudy (anonymous):

\[x=2 \pi+\frac{ \pi }{ 6 }=\frac{ 12 \pi+1 \pi }{ 6 }=?\]

OpenStudy (yesgeorgie8):

13pi/6

OpenStudy (anonymous):

yes you are correct. I suppose all are correct and you have to verify it.

OpenStudy (yesgeorgie8):

yes!!

OpenStudy (yesgeorgie8):

\[2cosx+1=0 \]

OpenStudy (anonymous):

1 . at x=pi/6\[2\sin x-1=2\sin \frac{ \pi }{ 6 }-1=2*\frac{ 1 }{ 2 }-1=1-1=0\] which is true.

OpenStudy (yesgeorgie8):

does that equation equal to 2pi/3 and 4pi/3

OpenStudy (anonymous):

which question you are talking about?

OpenStudy (yesgeorgie8):

2cosx+1=0

OpenStudy (anonymous):

it is 5th or 6th.

OpenStudy (yesgeorgie8):

solve equation: 2cosx+1=0, is the answer 2pi/3 and 4pi/3

OpenStudy (anonymous):

\[2\cos \frac{ 2 \pi }{ 3 }+1=2\cos \left( \pi-\frac{ \pi }{ 3 } \right)+1=-2\cos \frac{ \pi }{ 3 }+1=-2*\frac{ 1 }{ 2 }+1=0\] which is true. also it is true for 4 pi/3

OpenStudy (yesgeorgie8):

ok :)

OpenStudy (anonymous):

at x=2pi/3 \[2\cos ^2x-\cos x-1=2\cos ^2\frac{ 2\pi }{ 3 }-\cos \frac{ 2 \pi }{ 3 }-1=2\cos ^2\left( \pi-\frac{ \pi }{ 3 } \right)-\cos \left( \pi-\frac{ \pi }{ 3 } \right)-1\] \[=2\left[ -\cos \frac{ \pi }{ 3 } \right]^2-\left[ -\cos \frac{ \pi }{ 3 } \right]-1\] \[=2\cos ^2\frac{ \pi }{ 3 }+\cos \frac{ \pi }{ 3 }-1\] \[=2\left( \frac{ 1 }{ 2 } \right)^2+\frac{ 1 }{ 2 }-1\] \[=\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }-1=1-1=0\] which is true.